OpenStudy (anonymous):

At the local warehouse club where Billy buys supplies for his concession stand, a case of energy drinks cost twice as much as a case of water and a case of chips together. If a case of energy drinks costs $18 and a case of water costs $5, what is the cost of a case of chips?

OpenStudy (anonymous):

c = chips w = water e = energy drink 2w + c = e e = 18 w = 5 2(5) + c = 18 10 + c = 18 [subtract 10 from both sides] c = 8